# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 1. Introduction
In Section 2, we give formal definitions of classical and entangled CSPs, and we list some basic results about quantum homomorphisms. Section 3 presents the main definition of this work—quantum polymorphims—as well as their use to characterise the existence of oracular commutativity gadgets for entangled CSPs. In Section 4 we show that the quantum polymorphisms of cliques are non-contextual, which gives a different proof of the undecidability of the corresponding entangled CSPs due to [ji2013binary, Zeman]. In section 5, we prove a general result on the consequence of contextuality in quantum homomorphisms, by showing that it causes the emergence of a specific non-orthogonality pattern in the underlying structures. We make use of this pattern in section 6, to prove that odd cycles admit commutativity gadgets and, thus, the corresponding entangled CSP is undecidable. In Section 7, we address the non-oracular case, by showing that a similar characterisation to the one in Section 3 captures commutativity gadgets in the non-oracular setting. In addition, we establish therein a quantum version of Galois connection for the non-oracular quantum analogue of pp-definitions. Finally, in Section 8, we investigate the existence of commutativity gadgets for structures with a 2-element domain. We obtain that structures that do not admit the ternary majority as a polymorphism must have a commutativity gadget, and that apart from trivial cases, structures that do have the majority operation as a polymorphism do not have a commutativity gadget.
2. Preliminaries
2.1. Relational structures
@@
For $\ell\in\mathbb{N}$, $\ell$] is the set $\{1,2,\dots,\ell\}$. A signature $\sigma$ is a finite set of relation symbols $R$, each with its arity $\operatorname{ar}(R)\in\mathbb{N}$. A relational structure $\mathbb{X}$ with signature $\sigma$ (in short, a $\sigma$-structure or simply a structure when $\sigma$ is implicit) consists of a finite set $X$ and a relation $R^{\mathbb{X}}\subseteq X^{\operatorname{ar}(R)}$ for each symbol $R\in\sigma$. All relational structures in this work shall be implicitly assumed to be finite. A (classical) homomorphism $f:\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ the the entries of $\mathbf{x}$). We denote the existence of a classical homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ by the notation $\mathbb{X}\to\mathbb{Y}$. Given a $\sigma$-structure $\mathbb{A}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq A^{r}$, we define the structure $\mathbb{A}_{S}$ as follows. The signature of $\mathbb{A}_{S}$ contains a unique $r$-ary symbol $R$ (which does not belong to $\sigma$), and the interpretation of $R$ in $\mathbb{A}_{S}$ is $R^{\mathbb{A}_{S}}=S$. For typographical reasons, we shall sometimes denote $\mathbb{A}_{S}$ by $(A;S)$. For a digraph $\mathbb{X}$, we let $\mathbb{X}^{u}$ be the undirected graph on the same vertex set of $\mathbb{X}$, having an undirected edge $\{x,y\}$ whenever $(x,y)$ or $(y,x)$ form an edge in $\mathbb{X}$. We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless directed graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing a pair $(x,y)$ if, and only if, $x$ and $y$ appear together in some tuple in some relation of $\mathbb{X}$. Given a $\sigma$-structure $\mathbb{Y}$, the constraint satisfaction problem parameterised by $\mathbb{Y}$ (in short, $\operatorname{CSP}(\mathbb{Y})$) is the following computational problem: [^1] This is the decision version of $\operatorname{CSP}(\mathbb{Y})$. In the search version, the problem is to find an explicit homomorphism $f:\mathbb{X}\to\mathbb{Y}$ for any input $\mathbb{X}$ such that $\mathbb{X}\to\mathbb{Y}$. The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [bulatov2005classifying]. Given as input a $\sigma$-structure $\mathbb{X}$, output $\mathrm{YES}$ if $\mathbb{X}\to\mathbb{Y}$, and $\mathrm{NO}$ if $\mathbb{X}\not\to\mathbb{Y}$. For $\ell\in\mathbb{N}$, we denote by $\mathbb{Y}^{\ell}$ the $\ell$-th direct power of $\mathbb{Y}$; i.e., $\mathbb{Y}^{\ell}$ is the $\sigma$-structure whose domain is $Y^{\ell}$ and whose relations are defined as follows: Given $R\in\sigma$ and any $\ell\times\operatorname{ar}(R)$ matrix $M$ whose rows are tuples in $R^{\mathbb{Y}}$, the columns of $M$ form a tuple in $R^{\mathbb{Y}^{\ell}}$. A polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$. By $\operatorname{Pol}(\mathbb{Y})$ (resp. $\operatorname{Pol}^{(\ell)}(\mathbb{Y})$) we denote the set of all polymorphisms (resp. all polymorphisms of arity $\ell$) of $\mathbb{Y}$. The core of a $\sigma$-structure $\mathbb{X}$—in symbols, $\operatorname{core}(\mathbb{X})$—is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$. It is easy to check that $\operatorname{core}(\mathbb{X})$ is unique up to isomorphism, and any endomorphism is an automorphism (see e.g. [HellNesetrilCore]).
@@
The primitive positive (pp) fragment of first-order logic consists of the formulas that are built out of atomic formulas using conjunctions and existential quantifiers. If $\mathbb{A}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{A}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{A}$. If $R\subseteq A^{n}$ is a relation, we say that it is pp-definable over $\mathbb{A}$ if there exists a pp-formula defining it. A classical fact about pp-formulas is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{A}$, $(a_{1},\dots,a_{n})$ is in the relation defined by $\varphi$ in $\mathbb{A}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{A}$ that maps $x_{i}$ to $a_{i}$.
2.2. Quantum CSPs
@@
Let $H$ be a non-trivial finite-dimensional Hilbert space. We say that a linear endomorphism $p$ of $H$ is an orthogonal projector (or simply a projector) if it is self-adjoint and idempotent (i.e., if $p^{2}=p=p^{*}$). We denote by $\operatorname{id}_{H}$ (or simply by $\operatorname{id}$ when the Hilbert space is clear) the identity projector over $H$. A projective measurement (PVM) over a Hilbert space $H$ and a set $B$ and is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$. A quantum function from $A$ to $B$ is a family of PVMs $(Q_{a,b})_{b\in B}$, one PVM for each $a\in A$. We shall denote this by the notation $Q\colon A\to_{q}B$ or $Q\colon A\to_{H}B$ if we want to emphasize the Hilbert space. We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B,b^{\prime}\in B^{\prime}$. Given a quantum function $Q:A\to_{q}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$. We say that $Q$ is contextual if it is not non-contextual. Note that each classical function $f:A\to B$ can be seen as a non-contextual quantum function.
@@
Let $\mathbb{A},\mathbb{B}$ be relational structures with the same signature $\sigma$. A quantum homomorphism $Q\colon\mathbb{A}\to_{q}\mathbb{B}$, first defined at this level of generality in [abramsky2017quantum], is given by a non-trivial Hilbert space $H$ and a quantum function $Q\colon A\to_{q}B$ such that the following conditions hold:
@@
Note: Condition $($$\mathbf{QH_{2}}$)) means that $Q$ corresponds to perfect oracular correlations. In Section 7 we shall extend our results to the non-oracular case. We also point out that the notion of a quantum homomorphism can be phrased equivalently as a representation of a certain finitely generated *-algebra [MastelSlofstra2024].
Definition 1 (Entangled CSP).
@@
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{*}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{q}\mathbb{A}$.
2.3. Operations on quantum homomorphisms

### v2

@@ 2.8 Outlook
The remaining part of the paper is organised as follows. In Section 3, we formally define classical and quantum CSPs, and we give a list of some basic results about quantum homomorphisms and operations thereof that shall we useful throughout the paper. Section 4 presents the algebraic side of our result: we show in Section 4.1 that quantum polymorphisms are closed under minor operations and, thus, form an abstract minion, and in Section 4.2 we prove that minion homomorphism between quantum polymorphism minions determine complexity reductions between the corresponding quantum CSPs. We then turn to the relational side of classical/quantum bridge in Section 5, by formally describing q-definitions and q-constructions in Section 5.1, as well as establishing a Galois connection with quantum polymorphism. In Section 5.2 and Section 5.3, we show how the contextuality of quantum polymorphisms of a language fully captures commutativity gadgets, which translates into a fully polymorphic, concrete undecidability criterion for quantum CSPs (Theorem 7). This provably extends the known undecidability criteria in the literature on non-local games, as discussed in Section 5.4. In Section 6, we describe the quantum polymorphisms of cliques and obtain a simple proof of the undecidability of the corresponding CSPs, first established in [47, 33, 34]. Next, in Section 7, we describe non-orthogonality patterns for systems of orthogonal projectors that we call bifurcations, and we show that they occur in any contextual quantum homomorphism (or quantum polymorphism). We then specialise our analysis to particular classes of quantum CSPs. In particular, by using the polymorphic framework, contextuality bifurcations, and ad-hoc arguments based on the specific types of constraints, we establish in Section 8 and Section 9 the undecidability of the problems encoding quantum satisfiability of Siggers clauses and quantum homomorphisms to odd cycles, respectively. In Section 10, we consider the case of non-oracular quantum CSPs, we define the non-oracular version of quantum polymorphisms, and we show that they capture non-oracular commutativity gadgets in a similar way as in the oracular setting. Finally, Section 11 address quantum CSPs parameterised by Boolean languages. By describing the quantum polymorphisms of such languages, we obtain a full classification of the existence of commutativity gadgets, as well as a complete $\mathsf{P}$ vs. undecidable complexity dichotomy for Boolean quantum CSPs.
3 Preliminaries
In this section, we give formal definitions of most of the notions used throughout the paper, and we list a few simple facts on quantum homomorphisms and their operations.
3.1 Relational structures
@@
We let $\mathbb{N}=\{1,2,\dots\}$ denote the set of positive integers. For $\ell\in\mathbb{N}$, $\ell$] is the set $\{1,2,\dots,\ell\}$. A signature $\sigma$ is a finite set of relation symbols $R$, each with its arity $\operatorname{ar}(R)\in\mathbb{N}$. A relational structure $\mathbb{X}$ with signature $\sigma$ (in short, a $\sigma$-structure or simply a structure when $\sigma$ is implicit) consists of a finite set $X$ and a relation $R^{\mathbb{X}}\subseteq X^{\operatorname{ar}(R)}$ for each symbol $R\in\sigma$. All relational structures in this work shall be implicitly assumed to be finite. A (classical) homomorphism $f\colon\mathbb{X}\to\mathbb{Y}$ between two relational structures $\mathbb{X}$ and $\mathbb{Y}$ on a common signature $\sigma$ is a map from the domain $X$ of $\mathbb{X}$ to the domain $Y$ of $\mathbb{Y}$ that preserves all relations; i.e., $f(\mathbf{x})\in R^{\mathbb{Y}}$ for each symbol $R\in\sigma$ and each tuple $\mathbf{x}\in R^{\mathbb{X}}$ (where $f(\mathbf{x})$ is the entrywise application of $f$ to the entries of $\mathbf{x}$). We denote the existence of a classical homomorphism from $\mathbb{X}$ to $\mathbb{Y}$ by the notation $\mathbb{X}\to\mathbb{Y}$. Moreover, we let $\operatorname{Hom}(\mathbb{X},\mathbb{Y})$ be the set of all homomorphisms $f:\mathbb{X}\to\mathbb{Y}$. Given a $\sigma$-structure $\mathbb{X}$, a positive integer $r\in\mathbb{N}$, and a set $S\subseteq X^{r}$, we let $(X;S)$ be the structure with domain $X$ and a single relation $S$. We let the Gaifman graph of a relational structure $\mathbb{X}$ be the loopless undirected graph $\operatorname{Gaif}(\mathbb{X})$ on vertex set $V(\operatorname{Gaif}(\mathbb{X}))=X$ and edge set $E(\operatorname{Gaif}(\mathbb{X}))$ containing an edge $\{x,y\}$ if, and only if, $x$ and $y$ appear together in a tuple in some relation of $\mathbb{X}$.
Given a $\sigma$-structure $\mathbb{Y}$, the constraint satisfaction problem parameterised by $\mathbb{Y}$ (in short, $\operatorname{CSP}(\mathbb{Y})$) is the following computational problem: [^5] This is the decision version of $\operatorname{CSP}(\mathbb{Y})$. In the search version, the problem is to find an explicit homomorphism $f:\mathbb{X}\to\mathbb{Y}$ for any input $\mathbb{X}$ such that $\mathbb{X}\to\mathbb{Y}$. The two versions of $\operatorname{CSP}$ are equivalent up to polynomial-time reductions [22]. Given as input a $\sigma$-structure $\mathbb{X}$, output $\mathrm{YES}$ if $\mathbb{X}\to\mathbb{Y}$, and $\mathrm{NO}$ if $\mathbb{X}\not\to\mathbb{Y}$. For $\ell\in\mathbb{N}$, we denote by $\mathbb{Y}^{\ell}$ the $\ell$-th direct power of $\mathbb{Y}$; i.e., $\mathbb{Y}^{\ell}$ is the $\sigma$-structure whose domain is $Y^{\ell}$ and whose relations are defined as follows: given $R\in\sigma$ and any $\ell\times\operatorname{ar}(R)$ matrix $M$ whose rows are tuples in $R^{\mathbb{Y}}$, the columns of $M$ form a tuple in $R^{\mathbb{Y}^{\ell}}$. A (classical) polymorphism of $\mathbb{Y}$ (of arity $\ell$) is a homomorphism from $\mathbb{Y}^{\ell}$ to $\mathbb{Y}$. By $\operatorname{Pol}(\mathbb{Y})$ (resp. $\operatorname{Pol}^{(\ell)}(\mathbb{Y})$) we denote the set of all polymorphisms (resp. all polymorphisms of arity $\ell$) of $\mathbb{Y}$. The core of a $\sigma$-structure $\mathbb{X}$ is a minimal (under inclusion) induced substructure of $\mathbb{X}$ that is homomorphically equivalent to $\mathbb{X}$. It is easy to check that the core of a structure is unique up to isomorphism, and any endomorphism is an automorphism (see e.g. [43]).
@@
The primitive positive (pp) fragment of first-order logic consists of the formulae that are built out of atomic formulae using conjunctions and existential quantifiers. If $\mathbb{Y}$ is a relational structure and $\varphi(x_{1},\dots,x_{n})$ is a primitive positive formula in the signature of $\mathbb{Y}$ with free variables $x_{1},\dots,x_{n}$, then the relation defined by $\varphi$ is the set of $n$-tuples that satisfy the formula in $\mathbb{Y}$. If $R\subseteq Y^{n}$ is a relation, we say that it is pp-definable over $\mathbb{Y}$ if there exists a pp-formula defining it. A classical fact about pp-formulae is the following duality with gadgets: every pp-formula can be understood as a relational structure $\mathbb{X}_{\varphi}$ with distinguished vertices $x_{1},\dots,x_{n}$ such that for every structure $\mathbb{Y}$, $(y_{1},\dots,y_{n})$ is in the relation defined by $\varphi$ in $\mathbb{Y}$ if, and only if, there exists a homomorphism $\mathbb{X}_{\varphi}\to\mathbb{Y}$ that maps $x_{i}$ to $y_{i}$ for each $i\in[n]$. Suppose $\mathbb{X}$ and $\mathbb{Y}$ are relational structures with possible distinct domains and signatures. We say that $\mathbb{Y}$ has a pp-construction in $\mathbb{X}$ (or that $\mathbb{X}$ $p$-constructs $\mathbb{Y}$) if $\mathbb{Y}$ is homomorphically equivalent to a structure having domain $X^{d}$ for some $d\in\mathbb{N}$, and relations $R_{1},\dots,R_{k}$, where each $R_{i}\subseteq(X^{d})^{r_{i}}$ is pp-definable over $\mathbb{X}$, when seen as a relation $\tilde{R_{i}}\subseteq X^{dr_{i}}$ by “vectorising” tuples of tuples.
3.2 Quantum CSPs
@@
Let $H$ be a non-trivial finite-dimensional Hilbert space. We say that a linear endomorphism $p$ of $H$ is an orthogonal projector (or simply a projector) if it is self-adjoint and idempotent (i.e., if $p^{2}=p=p^{*}$). We denote by $\operatorname{id}_{H}$ (or simply by $\operatorname{id}$ when the Hilbert space is clear) the identity projector over $H$. A projective measurement (PVM) over a Hilbert space $H$ with outcome set $B$ is a family $(Q_{b})_{b\in B}$ of projectors on $H$ with the property that $\sum_{b\in B}Q_{b}=\operatorname{id}$. A quantum function from $A$ to $B$ is a family of PVMs $(Q_{a,b})_{b\in B}$, one PVM for each $a\in A$. We shall denote this by the notation $Q\colon A\to_{\operatorname{q}}B$ or $Q\colon A\to_{H}B$ if we want to emphasise the Hilbert space. We say that two PVMs $(Q_{b})_{b\in B}$ and $(Q^{\prime}_{b^{\prime}})_{b^{\prime}\in B^{\prime}}$ over the same Hilbert space commute if $Q_{b},Q^{\prime}_{b^{\prime}}]=0$ for each $b\in B$ and each $b^{\prime}\in B^{\prime}$. Given a quantum function $Q\colon A\to_{\operatorname{q}}B$ for two sets $A,B$, we say that $Q$ is non-contextual if $Q_{a,b},Q_{a^{\prime},b^{\prime}}]=0$ for each $a,a^{\prime}\in A$ and each $b,b^{\prime}\in B$. Otherwise, we say that $Q$ is contextual. Note that each classical function $f\colon A\to B$ can be seen as a non-contextual quantum function by setting $Q_{a,b}=\operatorname{id}$ if $f(a)=b$ and $Q_{a,b}=0$ otherwise.
@@
Let $\mathbb{A},\mathbb{B}$ be relational structures with the same signature $\sigma$. A quantum homomorphism $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$, first defined at this level of generality in [1], is given by a non-trivial finite-dimensional Hilbert space $H$ and a quantum function $Q\colon A\to_{\operatorname{q}}B$ such that the following conditions hold:
@@
It was proved in [1] that a quantum homomorphism $\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ exists if and only if the constraint-variable game associated with $\mathbb{A}$ and $\mathbb{B}$ admits a winning finite-dimensional tensor-product strategy. In the literature on non-local games, this is often expressed in terms of the existence of a perfect correlation in the set $C_{q}$ of quantum correlation matrices, see for example [67, 41, 46, 56]. Equivalently, the notion of a quantum homomorphism can be phrased as a representation of a certain finitely generated *-algebra [31, 64, 52, 60, 63]. Condition $($$\mathbf{QH_{2}}$)) is sometimes phrased as $Q$ expressing perfect oracular correlations; in Section 10 we shall extend our results to the non-oracular case.
Definition 14 (Quantum CSP).
@@
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP_{q}}(\mathbb{A})$ is the following computational problem: given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a quantum homomorphism $\mathbb{X}\to_{\operatorname{q}}\mathbb{A}$.
We will use the next observation on the length of walks in $\operatorname{Gaif}(\mathbb{A})$ and $\operatorname{Gaif}(\mathbb{B})$ for two structures $\mathbb{A}$ and $\mathbb{B}$ that are quantum homomorphic. This result is a slight modification of [59, Lemma 4.12] (which concerns the case where $\mathbb{A}$ and $\mathbb{B}$ are undirected graphs), and its proof closely follows that in [59].
Definition 15.
Let $\mathbb{A}$ be a relational structure with signature $\sigma$. We say that $\mathbb{A}$ is essentially binary if, for every $R\in\sigma$ and every $\mathbf{a}\in R^{\mathbb{A}}$, $\mathbf{a}$ has at most two distinct entries.
Proposition 16.
Let $Q\colon\mathbb{A}\to_{\operatorname{q}}\mathbb{B}$ be a quantum homomorphism between two structures $\mathbb{A}$ and $\mathbb{B}$ such that $\mathbb{A}$ is essentially binary. If $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$ are such that there exists an $\ell$-walk from $a$ to $a^{\prime}$ in $\operatorname{Gaif}(\mathbb{A})$ but there is no $\ell$-walk from $b$ to $b^{\prime}$ in $\operatorname{Gaif}(\mathbb{B})$, then $Q_{a,b}Q_{a^{\prime},b^{\prime}}=0$.
Proof.
We use induction over $\ell$. For $\ell=0$, the result is clear. Suppose $\ell=1$; i.e., $\{a,a^{\prime}\}\in E(\operatorname{Gaif}(\mathbb{A}))$ but $\{b,b^{\prime}\}\not\in E(\operatorname{Gaif}(\mathbb{B}))$. Assume, by contradiction, that $Q_{a,b}Q_{a^{\prime},b^{\prime}}\neq 0$. Since $\mathbb{A}$ is essentially binary, there exist some symbol $R\in\sigma$ and some tuple $\mathbf{z}=(z_{1},\dots,z_{r})\in R^{\mathbb{A}}$ such that $\{z_{1},\dots,z_{r}\}=\{a,a^{\prime}\}$. Consider now the tuple $\mathbf{w}=(w_{1},\dots,w_{r})\in B^{r}$ such that, for each $i\in[r]$, $w_{i}=b$ if $z_{i}=a$, and $w_{i}=b^{\prime}$ if $z_{i}=a^{\prime}$. Using $($$\mathbf{QH_{2}}$)) and the idempotency of projectors, we find that
$\prod_{i\in[r]}Q_{z_{i},w_{i}}=Q_{a,b}Q_{a^{\prime},b^{\prime}}\neq 0.$
We then deduce from $($$\mathbf{QH_{1}}$)) that $\mathbf{w}\in R^{\mathbb{B}}$, so $\{b,b^{\prime}\}\in E(\operatorname{Gaif}(\mathbb{B}))$; a contradiction. Take now $\ell\geq 2$, and suppose the result is true for all smaller values of $\ell$. Let $a_{0},a_{1},\dots,a_{\ell}$ be an $\ell$-walk in $\operatorname{Gaif}(\mathbb{A})$ such that $a_{0}=a$ and $a_{\ell}=a^{\prime}$. We have
$Q_{a,b}Q_{a^{\prime},b^{\prime}}=\sum_{z\in B}Q_{a,b}Q_{a_{\ell-1},z}Q_{a^{\prime},b^{\prime}}.$
Applying the inductive hypothesis to the $(\ell-1)$-walk $a_{0},\dots,a_{\ell-1}$ and the $1$-walk $a_{\ell-1},a_{\ell}$ in $\operatorname{Gaif}(\mathbb{A})$, we deduce that, in order for $z$ to yield a nonzero contribution in the summation above, $z$ must be connected to $b$ via an $(\ell-1)$-walk in $\operatorname{Gaif}(\mathbb{B})$, and to $b^{\prime}$ via a $1$-walk in $\operatorname{Gaif}(\mathbb{B})$. Concatenating the two walks yields an $\ell$-walk in $\operatorname{Gaif}(\mathbb{B})$ connecting $b$ to $b^{\prime}$. Hence, if no such walk exists, the product $Q_{a,b}Q_{a^{\prime},b^{\prime}}$ must be zero. ∎
3.3 Operations on quantum homomorphisms